OpenStudy (cenaida):

Solve the following system of equations using elimination: 3x - 4y = 8 10x + 5y = 45 A. (-4, -1) B. (3, 3) C. (4, 1) D. (8, 4)

OpenStudy (john_es):

Multiply the first equation by 10 and the second by -3, \[\begin{cases}30x-40y=80\\ -30x-15y=-135\end{cases}\]Now, sum the equations, \[-55y=-55\Rightarrow y=1\]Now, you can use again elimination, or simply use the first equation and substitute the y value obtained. Then solve for x, and you will obtain x=4. So the solution is (4,1).

OpenStudy (cenaida):

Thank You so much! John

OpenStudy (john_es):

;)